# First-Principles Framework for the Prediction of Intersystem Crossing Rates in Spin Defects: The Role of Electron Correlation

### v1

@@ Abstract
Our QDET calculations predict a $\Delta$ value of 0.414 eV, which exceeds the upper bound of $\Delta$ shown in Fig. 4(a), resulting in a $\Gamma_{A_{1}}$ rate of $66\pm 13$ MHz. This rate is slightly lower than the measured value of $100.5\pm 3.8$ MHz [48]. We also compare previously reported $\Delta$ values from various theoretical approaches [35, 55, 56, 57, 28, 29, 58] and compute the corresponding $\Gamma_{A_{1}}$ rates using $\lambda_{\perp,0}^{\textrm{QDET}}$ and $F_{A_{1}}(\Delta)$ from this work, as shown in Fig. 4(b). Notably, different theoretical approaches predict $\Delta$ in a broad range of $[0.04,0.51]$ eV, leading to $\Gamma_{A_{1}}$ between 20 MHz and 450 MHz. The values of $\Delta$ obtained from quantum chemistry approaches using carbon clusters are sensitive to both the level of theory and cluster size, resulting in $\Gamma_{A_{1}}$ variations by an order of magnitude. Overall, the comparison in Fig. 4(b) provides a valuable benchmark for quantitatively assessing the performance of various theoretical methods in predicting the energies of many-body states of spin defects. (See Sec. S9 [39] for a detailed comparison of ISC rates evaluated here and in previous studies [48, 49, 17, 29]).
@@
Figure 4: Intersystem crossing (ISC) rates and fluorescence lifetimes. (a) Computed ISC rate $\Gamma_{A_{1}}$ as a function of the energy gap ($\Delta$) between ${}^{3}\!E$ and ${}^{1}\!A_{1}$, with uncertainty (red shaded area) from the SOC parameter $\lambda_{\perp,0}^{\text{QDET}}$. The bounds $\Delta_{-}$ and $\Delta_{+}$ are determined by the intersection of the computed $\Gamma_{A_{1}}$ with experimental results [52]. $\Gamma_{A_{1}}$ from $\lambda_{\perp,0}^{\text{DFT}}$ is shown for comparison (dotted line). (b) $\Gamma_{A_{1}}$ computed using $\lambda_{\perp,0}^{\text{QDET}}$ and the vibrational overlap function $F_{A_{1}}(\Delta)$ from this work, with $\Delta$ values from various theoretical methods [35, 55, 56, 57, 28, 29, 58], including Franck-Condon shifts from TDDFT [35]. A visualization width of 0.02 eV is applied to each $\Delta$. Detailed data are provided in Table S5 [39]. (c) Computed fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ compared with experimental values from this work and previous studies by Batalov et al. [59], Robledo et al. [60], and Toyli et al. [61].
@@
Finally, we extend our computational framework for ISC rate calculations to finite temperatures ($T$) by summing contributions from different vibronic levels of ${}^{3}\!E$, weighted by the Boltzmann factor (see Sec. S8 [39]). We compute the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 K to 700 K, defined as $\tau(T)=1/\left[\Gamma_{\mathrm{Rad}}+\Gamma_{\mathrm{ISC}}(T)\right]$, where $\Gamma_{\mathrm{Rad}}=82.9$ MHz is the radiative decay rate of ${}^{3}\!E$ [52], and $\Gamma_{\mathrm{ISC}}(T)$ is the average ISC rate at a given $T$. To validate our calculations, we measured the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 K to 300 K in 50 K intervals. Combining these with previous measurements from Batalov et al. [59], Robledo et al. [60], and Toyli et al. [61], we obtain experimental data spanning 100–600 K. Details on the experimental setup and lifetime fitting are provided in Secs. S10 and S11 [39]. Despite moderate uncertainties in measurements and lifetime-fitting, our computed lifetimes agree well with experimental values below 600 K, as shown in Fig. 4(c). At temperatures above 600 K, a significant decrease in lifetime is observed experimentally, which may be attributed to additional decay pathways [61, 48] not yet included in our calculations, which will be explored in future work.
@@
Acknowledgments—The authors would like to thank Yuhang Ai for helpful discussions. The theoretical and computational work was primarily supported by the Midwest Integrated Center for Computational Materials (MICCoM) as part of the Computational Materials Sciences Program funded by the U.S. Department of Energy (Y.J., J.P., C.E. B.H., M.G., F.J.H., and G.G.). J.P. acknowledges the support from the Chicago Prize Postdoctoral Fellowship in Theoretical Quantum Science. Additional support for experimental validation includes the U.S. Department of Energy, Office of Science, Basic Energy Sciences Materials Sciences and Engineering Division (B.P. and D.D.A.), the SNU-Global Excellence Research Center establishment project and the NRF grant funded by the Korean government (D.D.O.) (MSIT) No. RS-2023-00258359, and from the AFOSR MURI under award No. FA9550-23-1-0330 (M.M.M. and C.B.). This research used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under contract No. DE-AC02-05CH11231 using NERSC Award No. ALCC-ERCAP0025950, and resources of the University of Chicago Research Computing Center.

### v2

@@ Abstract
Our QDET calculations predict a $\Delta$ value of 0.293 eV, which underestimates the lower bound of $\Delta$ shown in Fig. 4(a), resulting in a $\Gamma_{A_{1}}$ rate of $180\pm 34$ MHz. This rate is higher than the measured value of $100.5\pm 3.8$ MHz [48]. We also compare previously reported $\Delta$ values from various theoretical approaches [35, 55, 56, 57, 28, 29, 58] and compute the corresponding $\Gamma_{A_{1}}$ rates using $\lambda_{\perp,0}^{\textrm{QDET}}$ and $F_{A_{1}}(\Delta)$ from this work, as shown in Fig. 4(b). Notably, different theoretical approaches predict $\Delta$ in a broad range of $[0.04,0.51]$ eV, leading to $\Gamma_{A_{1}}$ between 20 and 450 MHz. The values of $\Delta$ obtained from quantum chemistry approaches using carbon clusters are sensitive to both the level of theory and cluster size, resulting in $\Gamma_{A_{1}}$ variations by an order of magnitude. Overall, the comparison in Fig. 4(b) provides a valuable benchmark for quantitatively assessing the performance of various theoretical methods in predicting the energies of many-body states of spin defects. (See Sec. S9 of SM [39] for a detailed comparison of ISC rates evaluated here and in previous studies [48, 49, 17, 29]).
@@
Figure 4: Intersystem crossing (ISC) rates and fluorescence lifetimes. (a) Computed ISC rate $\Gamma_{A_{1}}$ as a function of the energy gap ($\Delta$) between ${}^{3}\!E$ and ${}^{1}\!A_{1}$, with uncertainty (red shaded area) from the SOC parameter $\lambda_{\perp,0}^{\text{QDET}}$. The bounds $\Delta_{-}$ and $\Delta_{+}$ are determined by the intersection of the computed $\Gamma_{A_{1}}$ with experimental results [52]. $\Gamma_{A_{1}}$ from $\lambda_{\perp,0}^{\text{DFT}}$ is shown for comparison (dotted line). (b) $\Gamma_{A_{1}}$ computed using $\lambda_{\perp,0}^{\text{QDET}}$ and the vibrational overlap function $F_{A_{1}}(\Delta)$ from this work, with $\Delta$ values from various theoretical methods [35, 55, 56, 57, 28, 29, 58], including Franck-Condon shifts from TDDFT [35]. A visualization width of 0.02 eV is applied to each $\Delta$. Detailed data are provided in Table S5 of SM [39]. (c) Computed fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ compared with experimental values from this work and previous studies by Batalov et al. [59], Robledo et al. [60], and Toyli et al. [61].
@@
Finally, we extend our computational framework for ISC rate calculations to finite temperatures ($T$) by summing contributions from different vibronic levels of ${}^{3}\!E$, weighted by the Boltzmann factor (see Sec. S8 of SM [39]). This approach effectively accounts for the phonon-induced orbital averaging effect and yields the averaged ISC rate from the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ [48]. We compute the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 to 700 K, defined as $\tau(T)=1/\left[\Gamma_{\mathrm{Rad}}+\Gamma_{\mathrm{ISC}}(T)\right]$, where $\Gamma_{\mathrm{Rad}}=82.9$ MHz is the radiative decay rate of ${}^{3}\!E$ [52], and $\Gamma_{\mathrm{ISC}}(T)$ is the average ISC rate at a given $T$. To validate our calculations, we measured the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 to 300 K in 50 K intervals. Combining these with previous measurements from Batalov et al. [59], Robledo et al. [60], and Toyli et al. [61], we obtain experimental data spanning 100–600 K. Details on the experimental setup and lifetime fitting are provided in Secs. S10 and S11 of SM [39]. Despite moderate uncertainties in measurements and lifetime fitting, our computed lifetimes agree well with experimental values below 600 K, as shown in Fig. 4(c). At temperatures above 600 K, a significant decrease in lifetime is observed experimentally, which may be attributed to additional decay pathways [61, 48] not yet included in our calculations, which will be explored in future work.
@@
Acknowledgments—The authors would like to thank Siyuan Chen, Victor Wen-zhe Yu, and Yuhang Ai for helpful discussions. The theoretical and computational work was primarily supported by the Midwest Integrated Center for Computational Materials (MICCoM) as part of the Computational Materials Sciences Program funded by the U.S. Department of Energy (Y.J., J.P., C.E., B.H., M.G., F.J.H., and G.G.). J.P. acknowledges the support from the Chicago Prize Postdoctoral Fellowship in Theoretical Quantum Science. Additional support for experimental validation includes the U.S. Department of Energy, Office of Science, Basic Energy Sciences Materials Sciences and Engineering Division (B.P. and D.D.A.), the SNU-Global Excellence Research Center establishment project, and the NRF grant funded by the Korean government (D.D.O.) (MSIT) No. RS-2023-00258359, and from the AFOSR MURI under award No. FA9550-23-1-0330 (M.M.M. and C.B.). This research used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under contract No. DE-AC02-05CH11231 using NERSC Award No. ALCC-ERCAP0025950, and resources of the University of Chicago Research Computing Center.
Data availability—The data that support the findings of this Letter are openly available [66].